OpenStudy (anonymous):

Help Please!!!!!!!!!!! Ill Give Medal!!!!!!!! solve. z-5/7 = -1 z=???

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

\[\frac{ z-5 }{ 7 }= -1 \]

OpenStudy (anonymous):

:) kk thanks :)

OpenStudy (anonymous):

i think its z=-2

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

z=-2/7 :)

OpenStudy (anonymous):

:) thanks can you help me with some more??? :)

OpenStudy (anonymous):

ye why not

OpenStudy (anonymous):

Solve. 18 = – 2m + 6 – m m =

OpenStudy (anonymous):

m=-4

OpenStudy (anonymous):

since you've got 18=6-3m move 6 to the left becomes - so 12=-3m so m=-4

OpenStudy (anonymous):

isn't me whose suppose to take the medal?